OpenStudy (anonymous):

Solve for x: 3 over 4 x + 5 over 8 = 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (anonymous):

@jigglypuff314 @tHe_FiZiCx99 @thomaster

jigglypuff314 (jigglypuff314):

you can multiply every value by 8 to get rid of fractions for now so 3/4 x + 5/8 = 4x *8 *8 *8 to get 6x + 5 = 32x now does that look easier to solve? :)

OpenStudy (anonymous):

um can you help me

jigglypuff314 (jigglypuff314):

walk you through it? ok :) 6x + 5 = 32x you want to get all the values with x in them to one side -6x -6x so subtract both sides by 6x 5 = 26x but now there is a 26 attached to the x that you want to move to the other side it is attached to the x by multiplication so use division to move it divide both sides by 26 what would you get? :)

OpenStudy (anonymous):

0.193

jigglypuff314 (jigglypuff314):

yes :) but I think they want you to keep it as a fraction?

OpenStudy (anonymous):

so it would be?

OpenStudy (anonymous):

What? as a fraction can you turn it for me plz its really late

jigglypuff314 (jigglypuff314):

the fraction 5 divided by 26 :)

OpenStudy (anonymous):

thx

jigglypuff314 (jigglypuff314):

.-. you're welcome